OpenStudy (anonymous):

what is the length of the arc on a circle with radius 16ft intecepted by a 60 degree angel? round to the hundredth place pi=3.14

OpenStudy (michele_laino):

hint: we can write this proportion: \[\Large 360:\left( {2 \cdot 3.14 \cdot 16} \right) = 60:l\]